If the midpoint of a line segment is (3,4) amd one endpoint is (0,-6) what is the other end point? This question is not from a math book but from the 2007 Cedar Ridge Middle School Math Tournament 7th grade exam.
Answer by scott8148(6628)

the other endpoint is as far from the midpoint as the midpoint is from the 1st endpoint

if (x,y) is the other endpoint, then x-3=3-0 and y-4=4-(-6)

so x=6 and y=14